# Faithful Logic Embeddings in HOL – A recipe to have it all: deep and shallow, automated and interactive, heavy and light, proofs and counterexamples, meta and object level

### v1

@@ Abstract
Deep and shallow embeddings of non-classical logics in classical higher-order logic have been explored, implemented, and used in various automated reasoning tools in recent years. This paper presents a recipe for the simultaneous deployment of different forms of deep and shallow embeddings in classical higher-order logic, enabling not only flexible interactive and automated theorem proving and counterexample finding at meta and object level, but also automated faithfulness proofs between the logic embeddings. The approach, which is fruitful for logic education, research and application, is deliberately illustrated here using simple propositional modal logic. However, the work presented is conceptual in nature and not limited to such a simple logic context.
@@ 1 Motivation and Introduction
Deep embeddings of logics, or more generally of domain-specific languages, in a suitable metalogic, such as the classical higher-order logic (HOL) [23, 4], are typically based on explicitly introduced abstract data types that essentially axiomatize the inductively defined character of the new language. Shallow embeddings, on the other hand, identify the embedded logics or languages as fragments of the metalogic and they typically require few or no axiom postulates. Since deep and shallow embeddings logics in HOL effectively and elegantly support, for example, interactive and automated reasoning with and about such logics in already existing theorem proving systems for the chosen metalogic, they have received increased attention in recent years. This becomes even more relevant if, for example, quantified non-classical logics are embedded and mechanized this way, since many such logics have relevant applications, but often only a few or none direct prover implementations exist. This is the case, for example, for deontic logics which e.g. have topical applications in normative reasoning and AI ethics.
@@
Against this background, the motivation and idea of this paper is multifaceted. First, this paper presents a technique for inherently linking deep and shallow logic embeddings in metalogic HOL, with the effect that computer-supported faithfulness proofs between such embeddings are not only enabled, but may now even be automated in HOL; this is, to the best of the authors’ knowledge, a novel result. Second, the paper discusses different forms of shallow embeddings, which differ in the degree of semantic dependencies that are explicitly encoded and maintained, while the remaining dependencies are shared with the metalogic HOL. While this paper contrasts heavyweight, maximal, and lightweight, minimal, shallow embeddings as extremes, it should become clear that there is a whole spectrum between these extremes that typically exists. Illustrating this aspect is a second novel contribution, which is useful also for obtaining a better understanding and categorization of existing prior work. Third, the paper points to different possibilities for logic education and for research about and with logic embeddings within modern proof assistants and theorem provers for HOL. These range from metalogical studies to object logic applications using interactive proof and disproof, as well as automated theorem proving and counterexample finding, while allowing precise control over the particular level, e.g. object- or metalevel, at which reasoning takes place. Fourth, the paper conjectures that minimal shallow embeddings seem to be better suited for certain reasoning tasks than their maximal or deep embedding counterparts and it provides evidence for this. Fifth, using the Löb axiom of provability logic, the paper briefly sketches how research on the detection of semantic correspondences could possibly be supported by the presented technique in the future.
As an exemplary non-classical logic to illustrate the above aspects, simple propositional modal logic (PML) [29, 15, 49] is deliberately chosen here, since it may be well known to the wider audience that this paper intends to reach.
@@
Moreover, Isabelle/HOL [44] was chosen as the host proof assistant system for the studies presented here. Conceptually, however, other theorem provers for HOL could have been used instead. Reasons for using Isabelle/HOL include (i) its currently unmatched support for proof automation with sledgehammer [16] and various other tools, (ii) the availability of the model finder nitpick [17] for generating counterexamples, and (iii) its powerful user interface, which allows for user-friendly -quality representations. In principle, this paper could have taken advantage of Isabelle/HOL’s special support for computer-verifiable paper writing to generate a document directly from the Isabelle/HOL sources, as is the case, for example, with all the articles published in the Archive of Formal Proofs. However, the author has decided against this idea and in favor of screenshots of the source code, as this allows for more compact representations that saved space and, in the author’s opinion, are even somewhat more didactically useful due to visual highlighting and indenting, putting a better focus on relevant information.
@@
Organization of the paper: §2 and §3 briefly introduce PML and HOL, before §4-6 present deep, maximally shallow, and minimally shallow embeddings of PML in HOL. These are connected by automated faithfulness proofs given in §7. Various experiments with the presented embeddings are presented in §8, before §9 discusses related work and §10 concludes the paper. Knowledgeable readers may simply skip §2 and §3, which are included to ensure self-containedness.
2 Popositional modal logic (PML)
@@
A semantics for PML, called relational semantics or possible-world semantics, determines the truth of a formula $\varphi$ relative to a so-called possible world $w$ and a relational model. A relational model is a tuple $\langle\mathcal{W},\mathcal{R},\mathcal{V}\rangle$, where $\mathcal{W}$ is a set of possible worlds (called the universe), $\mathcal{R}$ is a binary relation on $W$ (called the accessibility relation), $\mathcal{V}$ is a valuation function that assigns a truth value to each symbol $a\in\mathcal{S}$ in each world $w\in\mathcal{W}$, where $\mathcal{S}$ is the signature of propositional symbols. Formally, the truth of a formula $\varphi$ with respect to a given model $\langle\mathcal{W},\mathcal{R},\mathcal{V}\rangle$ and a possible world $w\in\mathcal{W}$ is defined as follows:
@@
$\langle\mathcal{W},\mathcal{R},\mathcal{V}\rangle,w$ $\models\varphi\supset\psi\text{\ iff\ }\langle\mathcal{W},\mathcal{R},\mathcal{V}\rangle,w\models\varphi\text{\ implies\ }\langle\mathcal{W},\mathcal{R},\mathcal{V}\rangle,w$
$\langle\mathcal{W},\mathcal{R},\mathcal{V}\rangle,w$ $\models\Box\varphi\text{\ iff\ for\ all\ }v\text{\ with\ }Rwv\text{\ we have\ }\langle\mathcal{W},\mathcal{R},\mathcal{V}\rangle,v\models\varphi$
Note that in this recursive definition of truth the only parameter that may be recursively modified is the possible world $w$ in the $\Box$-clause, which states that a formula $\varphi$ is “necessary” a world $w$ iff $\varphi$ is true at all reachable worlds $v$ from $w$. Based on the above definitions a notion of validity can be defined: A formula $\varphi$ is valid (in so called base modal logic K) iff $\langle\mathcal{W},\mathcal{R},\mathcal{V}\rangle,w\models\varphi$ for all $\mathcal{W}$, $\mathcal{R}$, $\mathcal{V}$, and $w\in\mathcal{W}$.
@@
A proof system for modal logic K is obtained by adding the schematic necessitation rule, stating that “if $\varphi$ is a theorem of K, then so is $\Box\varphi$” and the distribution axiom schema $\Box(\varphi\supset\psi)\supset(\Box\varphi\supset\Box\psi)$ to the principles of classical propositional logic, which in turn can be captured by a simple Hilbert calculus [56, 50] consisting of the modes ponens rule, “if $\varphi$ and $\varphi\supset\psi$ are theorems (of K), then so is $\psi$”, and the three axiom schemata H1: $\varphi\supset(\psi\supset\varphi)$, H2: $(\varphi\supset(\psi\supset\gamma))\supset((\varphi\supset\psi)\supset(\varphi\supset\gamma))$ and H3: ($(\neg\varphi\supset\neg\psi)\supset(\psi\supset\varphi)$).
@@ 3 Classical higher-order logic (HOL)
[…]
@@ 4 Deep embedding of PML in HOL
This section presents a deep embedding of PML in HOL. Both the PML language and its semantics are formalized very closely following the definitions given in §2. This formalization, and all others presented in this paper, are based on some uniformly assumed conventions, resp. preliminaries. These preliminaries are provided in a file PMLinHOL_preliminaries.thy; see Appendix 0.A. [^1] A preprint of this paper is available online at arXiv, and the source package provided there also contains all the Isabelle/HOL sources of the presented work. Relevant at this point is that a type $\mathcal{S}$ is introduced to represent the signature $\mathcal{S}$ of PML, i.e. a (finite or infinite) set of propositional logic symbols; type w is the type of possible worlds, $\mathcal{W}:=\texttt{w}{\Rightarrow}\texttt{bool}$ is the predicate type of possible worlds (we regard predicates of this type as characteristic functions for the associated set of possible worlds), $\mathcal{R}:=\texttt{w}{\Rightarrow}\texttt{w}{\Rightarrow}\texttt{bool}$ is the type of accessibility relations between possible worlds, and $\mathcal{V}:=\mathcal{S}{\Rightarrow}\texttt{w}{\Rightarrow}\texttt{bool}$ is the type of evaluation functions. $\forall w{:}W.\,\varphi$ is a $W$-predicate guarded quantifier that unfolds into $\forall w.\,W\,w\longrightarrow\varphi$ ($W$ is of type $\mathcal{W}$). An important remark concerns the sharing of the signature type $\mathcal{S}$ between all embeddings introduced in this paper. This sharing is essential for enabling the faithfulness proofs between the embeddings as presented in §7.
@@
The deep embedding of PML in HOL is defined in file PMLinHOL_deep.thy and presented at top of Fig. 1, in which comments are encapsulated in orange delimiters <>{-}{<}\ {>}- < >. After importing PMLinHOL_preliminaries.thy, this file starts out with defining an abstract datatype [^2] Such a datatype declaration corresponds to, and unfolds into, a rather large set of axioms in HOL that could be postulated instead. Using the Isabelle/HOL command print_theorems, this set can be inspected by the user. introducing the inductively defined logic language PML, which introduces atomic formulas $a^{d}$ (for $a\in\mathcal{S}$), negated formulas $\neg^{d}\varphi$, implication formulas $\varphi\vee^{d}\psi$, and boxed formulas $\Box^{d}\varphi$ (see line 4). The superscript s is used to mark all deeply embedded connectives so that they can be properly distinguished from their counterparts defined in the following sections. In these sections, the alternative superscripts s (maximal shallow) and m (minimal shallow) are used accordingly.
@@
In accordance with §2 the further logical connectives $\vee^{d}$, $\wedge^{d}$, $\Diamond^{d}$, $\top^{d}$, and $\bot^{d}$ are introduced as defined terms (lines 6-10). [^3] Abbreviations could be used instead. They are always unfolded implicitly, while definition unfolding is an explicit operation, leading to more options for control. In any case, readers new to Isabelle/HOL can simply focus on the indented definition equations here and in other figures, at least at first reading. Then the primitive recursive definition of truth in a world $w$ for a given model $\langle{W},{R},{V}\rangle$, termed RelativeTruthD and noted as $\langle{W},{R},{V}\rangle,w\models^{d}\varphi$, is given (lines 12-16). This definition follows one to one, even in presentation style, the definition introduced in §2. Based on this, the validity $\models^{d}\varphi$ of a deeply embedded formula $\varphi$ is defined as relative truth wrt. all sets of worlds $W$, accessibility relations $R$, valuation functions $V$, and worlds $w\in W$ (line 18). Finally, in order to support a convenient, simultaneous unfolding of the introduced definitions, a bag DefD is defined (in lines 20-22), which bundles them and adds them also to the simplifier simp of Isabelle/HOL.
@@ 5 Maximal shallow embedding of PML in HOL
Truth in a world $w$ for a given model $\langle{W},{R},{V}\rangle$, called RelativeTruthS and noted as $\langle{W},{R},{V}\rangle,w\models^{s}\varphi$, is then defined (see line 17) as the application of $\varphi$ to ${W}$, ${R}$, ${V}$ and $w$. Analogous to the deep embedding from above, validity $\models^{s}\varphi$ of an embedded formula $\varphi$ is defined as relative truth wrt. all sets of worlds $W$, accessibility relations $R$, valuation functions $V$, and worlds $w\in W$ (line 19). Also analogous to above, a bag DefS is defined (lines 21-23), which conveniently bundles all the definitions given in this file and adds them to the simplifier. Note that not a single axiom is being introduced in this maximal shallow embedding, with the effect, that PML is being characterized here as a proper, albeit quite heavyeight, fragment of the language of HOL.
@@ 6 Minimal shallow embedding of PML in HOL
To implement this idea, ${W}$ is now implicitly identified with type w of HOL at the metalevel, and the constant symbols R and V are introduced in metalogic HOL to represent the semantic notions ${R}$ and ${V}$ of PML (see line 5). Most importantly, the modified HOL type $\sigma:=\texttt{w}{\Rightarrow}\texttt{bool}$ is now taken as the type of the embedded formulas $\varphi_{\sigma}$, indicating that PML formulas in this minimal embedding now depend only on possible worlds. [^5] Note that this is also the core idea of the well-known standard translation of PML into first-order logic [15], which, in fact, is implemented here by exploiting $\lambda$-conversion and compositionality directly in the metalogical HOL without the need for explicit recursive definitions. Hence, instead of maintaining dependencies on ${W}$, ${R}$, ${V}$, and $w$, the definitions of the logical connectives $a^{m}$, $\neg^{m}$, $\supset^{m}$, and $\Box^{m}$ now only explicitly maintain a dependency on worlds $w$ (lines 8-11). Their definitions are thus, where V and R are the above-mentioned constant symbols in metalogic HOL: $a^{s}:=\lambda w.\,\texttt{V}\,a\,w$, $\neg^{m}\varphi:=\lambda w.\,\neg(\varphi\,w)$ and $\varphi\supset^{m}\psi:=\lambda w.\,\varphi\,w\longrightarrow\psi\,w$, and $\Box^{m}\varphi:=\lambda w.\,\forall v.\,\texttt{R}\,w\,v\longrightarrow\varphi\,v$. The definition of the other logical connectives $\vee^{m}$, $\wedge^{m}$, $\Diamond^{m}$, $\top^{m}$, and $\bot^{m}$ (in lines 13-17) is aanalogous to before. Truth in a world $w$, for a given model $\langle{W},{R},{V}\rangle$ maintained at the meta-lavel HOL, called RelativeTruthM and noted as $w\models^{m}\varphi$, is then defined (see line 19) as the application of the embedded formula $\varphi$ to $w$. Consequently, validity $\models^{m}\varphi$ of an embedded formula $\varphi$ is defined as relative truth wrt. to all worlds $w$ (line 21). The bag DefM defined (in lines 21-23) is analogous to DefD and DefS from before. Again, not a single axiom is introduced. So PML is once again characterized as a proper fragment of HOL, but now more lightweight.

### v2

# Faithful Logic Embeddings in HOL — Deep and Shallow

@@ Abstract
Deep and shallow embeddings of non-classical logics in classical higher-order logic have been explored, implemented, and used in various reasoning tools in recent years. This paper presents a method for the simultaneous deployment of deep and shallow embeddings of various degrees in classical higher-order logic. This enables flexible, interactive and automated theorem proving and counterexample finding at meta and object level, as well as automated faithfulness proofs between these logic embeddings. The method is beneficial for logic education, research and application and is illustrated here using a simple propositional modal logic. However, this approach is conceptual in nature and not limited to this simple logical context.
@@ 1 Motivation and Introduction
Deep embeddings of logics, or more generally of domain-specific languages, in a suitable metalogic, such as the classical higher-order logic (HOL) [27, 6], are typically based on explicitly introduced abstract data types that essentially axiomatize the inductively defined character of the new language. Shallow embeddings, on the other hand, identify the embedded logics or languages as fragments of the metalogic and they typically require few or no axiom postulates. [^1] Gibbons and Wu characterize the two notions as follows [34, p.2]: “[…] a deep embedding consists of a representation of the abstract syntax as an algebraic datatype, together with some functions that assign semantics to that syntax by traversing the algebraic datatype. A shallow embedding eschews the algebraic datatype, and hence the explicit representation of the abstract syntax of the language; instead, the language is defined directly in terms of its semantics.” Since deep and shallow embeddings logics in HOL effectively and elegantly support, for example, interactive and automated reasoning with and about such logics in already existing theorem proving systems for HOL, embeddings in HOL have received increased attention in recent years, see §9. Shallow embeddings are e.g. well suited to automate quantified non-classical logics (which means that they can be used in areas such as computational metaphysics [16]), and deep embeddings are e.g. well suited to reason about the syntactic structure of embedded logic.
@@
The contributions of this paper are manifold: First, a technique for the simultaneous provision of deep and shallow embeddings in metalogic HOL is presented, with the effect that computer-supported faithfulness proofs between the embeddings are not only enabled, but may — at least for simple propositional object logics —now even be fully automated in HOL; this is, to the best of the authors’ knowledge, a novel result. Second, this paper discusses shallow embedding alternatives differing in the degree of semantic dependencies that are explicitly encoded and maintained. More precisely, it contrasts heavyweight (maximal) and lightweight (minimal) shallow embeddings as extremes. Doing so it hints at the full spectrum that exists between these extremes, which is in fact relevant for practice. For example, while a minimal flat embedding for a multimodal logic only needs to explicitly encode the crucial semantic dependency on possible worlds [11], an additional dependency, namely on a now possibly dynamically changing set of considered possible worlds, becomes crucial if the logic is extended by a public announcement operator [13]. Third, the paper hints at opportunities for logic education, and for research, about and with logic embeddings within modern proof assistants and theorem provers for HOL. These range from metalogical studies to object logic applications using interactive proof and disproof, as well as automated theorem proving and counterexample finding, while allowing precise control over the particular level, e.g. object- or metalevel, at which reasoning takes place. Fourth, the paper presents some preliminary studies on whether shallow or deep embeddings are better suited for certain reasoning tasks. Fifth, using the Löb axiom of provability logic, the paper briefly sketches how research on the detection of semantic correspondences could possibly be supported by the presented technique in the future.
@@
Since this paper and associated material is intended to be used in logic education, simple propositional modal logic (PML) [33, 18, 55] has been chosen to illustrate the addressed ideas, and Isabelle/HOL [49] was chosen as the host proof assistant system for the experiments conducted. Reasons for using Isabelle/HOL include (i) its very good proof automation support with e.g. the sledgehammer [19] tool, (ii) the availability of the (counter-)model finder nitpick [20], and (iii) its powerful user interface, which allows for user-friendly -quality representations. In principle, this paper could have taken advantage of Isabelle/HOL’s special support for computer-verifiable paper writing to generate a document directly from the Isabelle/HOL sources, as is the case, for example, with all the articles published in the Archive of Formal Proofs. However, the author has decided against this idea and in favor of screenshots of the source code, as this allows for more compact representations that saved space and, in the author’s opinion, are even somewhat more didactically useful due to visual highlighting and indenting, putting a better focus on relevant information.
@@
Organization of the paper: §2 and §3 briefly introduce PML and HOL, before §4-6 present deep, maximally shallow, and minimally shallow embeddings of PML in HOL. These are connected by automated faithfulness proofs given in §7. Various experiments with the presented embeddings are presented in §8, before §9 discusses related work and §10 concludes the paper.
2 Propositional modal logic (PML)
@@
A semantics for PML, called relational semantics or possible-world semantics, determines the truth of a formula $\varphi$ relative to a so-called possible world $w$ and a relational model. A relational model is a tuple $\langle\mathcal{W},\mathcal{R},\mathcal{V}\rangle$, where $\mathcal{W}$ is a set of possible worlds (called the universe), $\mathcal{R}$ is a binary relation on $W$ (called the accessibility relation), $\mathcal{V}$ is a valuation function that assigns a truth value to each symbol $a\in\mathcal{S}$ in each world $w\in\mathcal{W}$, where $\mathcal{S}$ is the signature of propositional symbols. $\langle\mathcal{W},\mathcal{R}\rangle$ is called a frame. Formally, the truth of a formula $\varphi$ with respect to a given model $\langle\mathcal{W},\mathcal{R},\mathcal{V}\rangle$ and a possible world $w\in\mathcal{W}$ is defined as follows:
@@
$\langle\mathcal{W},\mathcal{R},\mathcal{V}\rangle,w$ $\models\varphi\supset\psi\text{\ iff\ }\langle\mathcal{W},\mathcal{R},\mathcal{V}\rangle,w\models\varphi\text{\ implies\ }\langle\mathcal{W},\mathcal{R},\mathcal{V}\rangle,w\models\psi$
$\langle\mathcal{W},\mathcal{R},\mathcal{V}\rangle,w$ $\models\Box\varphi\text{\ iff\ for\ all\ }v\text{\ with\ }\mathcal{R}wv\text{\ we have\ }\langle\mathcal{W},\mathcal{R},\mathcal{V}\rangle,v\models\varphi$
Note that in this recursive definition of truth the only parameter that may be recursively modified is the possible world $w$ in the $\Box$-clause, which states that a formula $\varphi$ is “necessary” in a world $w$ iff $\varphi$ is true at all reachable worlds $v$ from $w$. Based on the above definitions a notion of validity can be defined: A formula $\varphi$ is valid (in so called base modal logic K) iff $\langle\mathcal{W},\mathcal{R},\mathcal{V}\rangle,w\models\varphi$ for all $\mathcal{W}$, $\mathcal{R}$, $\mathcal{V}$, and $w\in\mathcal{W}$.
@@
A proof system for modal logic K is obtained by adding the schematic necessitation rule, stating that “if $\varphi$ is a theorem of K, then so is $\Box\varphi$” and the distribution axiom schema $\Box(\varphi\supset\psi)\supset(\Box\varphi\supset\Box\psi)$ to the principles of classical propositional logic. The latter can be captured by a simple Hilbert calculus [62, 56] consisting of the modes ponens rule, “if $\varphi$ and $\varphi\supset\psi$ are theorems (of K), then so is $\psi$”, and the three axiom schemata H1: $\varphi\supset(\psi\supset\varphi)$, H2: $(\varphi\supset(\psi\supset\gamma))\supset((\varphi\supset\psi)\supset(\varphi\supset\gamma))$ and H3: ($(\neg\varphi\supset\neg\psi)\supset(\psi\supset\varphi)$).
@@ 3 Classical higher-order logic (HOL)
$s_{o}$ is valid in $\mathcal{M}$ under assignment $g$, denoted as $\mathcal{M},g\models^{\texttt{HOL}}s_{o}$, iff $\llbracket s_{o}\rrbracket^{\mathcal{M},g}$$=T$ $s_{o}$ is valid in $\mathcal{M}$, denoted as $\mathcal{M}\models^{\texttt{HOL}}s_{o}$, iff $\mathcal{M},g\models^{\texttt{HOL}}s_{o}$ for all assignments $g$, and $s_{o}$ is valid, denoted as $\models^{\texttt{HOL}}s_{o}$, iff $s_{o}$ is valid in all Henkin models $\mathcal{M}$.
@@ 4 Deep embedding of PML in HOL
This section presents a deep embedding of PML in HOL. Both the PML language and its semantics are formalized very closely following the definitions given in §2. This formalization, and all others presented in this paper, are based on some uniformly assumed conventions, resp. preliminaries. These preliminaries are provided in a file PMLinHOL_preliminaries.thy; see Fig 1.
@@
Relevant at this point is that a type $\mathcal{S}$ is introduced to represent the signature $\mathcal{S}$ of PML, i.e. a (finite or infinite) set of propositional logic symbols; type w is the type of possible worlds, $\mathcal{W}:=\texttt{w}{\Rightarrow}\texttt{bool}$ is the predicate type of possible worlds (we regard predicates of this type as characteristic functions for the associated set of possible worlds), $\mathcal{R}:=\texttt{w}{\Rightarrow}\texttt{w}{\Rightarrow}\texttt{bool}$ is the type of accessibility relations between possible worlds, and $\mathcal{V}:=\mathcal{S}{\Rightarrow}\texttt{w}{\Rightarrow}\texttt{bool}$ is the type of evaluation functions. Well known predicates and functions on relations are introduced (in lines 11-19), and $\forall w{:}W.\,\varphi$ (introduced in lines 21-23) is a $W$-predicate guarded quantifier that unfolds into $\forall w.\,W\,w\longrightarrow\varphi$ ($W$ is of type $\mathcal{W}$). An important remark concerns the sharing of the signature type $\mathcal{S}$ between all embeddings introduced in this paper. This sharing is essential for enabling the faithfulness proofs between the embeddings as presented in §7.
@@
The deep embedding of PML in HOL is defined in file PMLinHOL_deep.thy and presented at top of Fig. 2, in which comments are encapsulated in orange delimiters <>{-}{<}\ {>}- < >. After importing PMLinHOL_preliminaries.thy, this file starts out with defining an abstract datatype [^2] Such a datatype declaration corresponds to, and unfolds into, a rather large set of axioms in HOL that could be postulated instead. Using the Isabelle/HOL command print_theorems, this set can be inspected by the user. introducing the inductively defined logic language PML, which introduces atomic formulas $a^{d}$ (for $a\in\mathcal{S}$), negated formulas $\neg^{d}\varphi$, implication formulas $\varphi\supset^{d}\psi$, and boxed formulas $\Box^{d}\varphi$ (see line 4). The superscript d is used to mark all deeply embedded connectives so that they can be properly distinguished from their counterparts defined in the following sections. In these sections, the alternative superscripts s (maximal shallow) and m (minimal shallow) are used accordingly.
@@
In accordance with §2 the further logical connectives $\vee^{d}$, $\wedge^{d}$, $\Diamond^{d}$, $\top^{d}$, and $\bot^{d}$ are introduced as defined terms (lines 6-10). [^3] Abbreviations could be used instead. They are always unfolded implicitly, while definition unfolding is an explicit operation, leading to more options for control. In any case, readers new to Isabelle/HOL can simply focus on the indented definition equations here and in other figures, at least at first reading. Then the primitive recursive definition of truth in a world $w$ for a given model $\langle{W},{R},{V}\rangle$, termed RelativeTruthD and noted as $\langle{W},{R},{V}\rangle,w\models^{d}\varphi$, is given (lines 12-16). This definition follows one to one, even in presentation style, the definition introduced in §2. Based on this, the validity $\models^{d}\varphi$ of a deeply embedded formula $\varphi$ is defined as relative truth w.r.t. all sets of worlds $W$, accessibility relations $R$, valuation functions $V$, and worlds $w\in W$ (line 18). Finally, in order to support a convenient, simultaneous unfolding of the introduced definitions, a bag DefD is defined (in lines 20-22), which bundles them and adds them also to the simplifier simp of Isabelle/HOL.
Figure 1: Preliminaries.
@@ 5 Maximal shallow embedding of PML in HOL
Truth in a world $w$ for a given model $\langle{W},{R},{V}\rangle$, called RelativeTruthS and noted as $\langle{W},{R},{V}\rangle,w\models^{s}\varphi$, is then defined (see line 17) as the application of $\varphi$ to ${W}$, ${R}$, ${V}$ and $w$. Analogous to the deep embedding from above, validity $\models^{s}\varphi$ of an embedded formula $\varphi$ is defined as relative truth w.r.t. all sets of worlds $W$, accessibility relations $R$, valuation functions $V$, and worlds $w\in W$ (line 19). Also analogous to above, a bag DefS is defined (lines 21-23), which conveniently bundles all the definitions given in this file and adds them to the simplifier. Note that not a single axiom is being introduced in this maximal shallow embedding, with the effect, that PML is being characterized here as a proper, albeit quite heavyeight, fragment of the language of HOL.
@@ 6 Minimal shallow embedding of PML in HOL
To implement this idea, ${W}$ is now implicitly identified with type w of HOL at the metalevel, and the constant symbols R and V are introduced in metalogic HOL to represent the semantic notions ${R}$ and ${V}$ of PML (see line 5). Most importantly, the modified HOL type $\sigma:=\texttt{w}{\Rightarrow}\texttt{bool}$ is now taken as the type of the embedded formulas $\varphi_{\sigma}$, indicating that PML formulas in this minimal embedding now depend only on possible worlds. [^5] Note that this is also the core idea of the well-known standard translation of PML into first-order logic [18], which, in fact, is implemented here by exploiting $\lambda$-conversion and compositionality directly in the metalogical HOL without the need for explicit recursive definitions. Hence, instead of maintaining dependencies on ${W}$, ${R}$, ${V}$, and $w$, the definitions of the logical connectives $a^{m}$, $\neg^{m}$, $\supset^{m}$, and $\Box^{m}$ now only explicitly maintain a dependency on worlds $w$ (lines 8-11). Their definitions are thus, where V and R are the above-mentioned constant symbols in metalogic HOL: $a^{s}:=\lambda w.\,\texttt{V}\,a\,w$, $\neg^{m}\varphi:=\lambda w.\,\neg(\varphi\,w)$ and $\varphi\supset^{m}\psi:=\lambda w.\,\varphi\,w\longrightarrow\psi\,w$, and $\Box^{m}\varphi:=\lambda w.\,\forall v.\,\texttt{R}\,w\,v\longrightarrow\varphi\,v$. The definition of the other logical connectives $\vee^{m}$, $\wedge^{m}$, $\Diamond^{m}$, $\top^{m}$, and $\bot^{m}$ (in lines 13-17) is analogous to before. Truth in a world $w$, for a given model $\langle{W},{R},{V}\rangle$ maintained at the meta-lavel HOL, called RelativeTruthM and noted as $w\models^{m}\varphi$, is then defined (see line 19) as the application of the embedded formula $\varphi$ to $w$. Consequently, validity $\models^{m}\varphi$ of an embedded formula $\varphi$ is defined as relative truth w.r.t. to all worlds $w$ (line 21). The bag DefM defined (in lines 21-23) is analogous to DefD and DefS from before. Again, not a single axiom is introduced. So PML is once again characterized as a proper fragment of HOL, but now more lightweight.